# Refactoring Codebases through Library Design

### v1

@@ 4 Librarian: Refactoring Code to Create Libraries
,$\mathcal{L}^{\star},\left\{\rho^{\star}_{n}\right\}=\operatorname*{arg\,min}_{\mathcal{L},\left\{\rho^{\prime}_{n}\right\}\in\textsc{Sample}\left(\left\{\rho_{n}\right\}\right)}\ell\left(\mathcal{L},\left\{\rho^{\prime}_{n}\right\}\right).$ (2)
4.1 Sample with clustering
Meaningful abstractions exist primarily among programs that share some functionality or underlying structure. We perform clustering on the input programs to make groups of programs that likely share structures that can be abstracted into general library functions. Most modern language models cannot be prompted with the entire collection of input programs– even long context models cannot process the entirety of e.g the Linux kernel, and even if they could, it is not clear that such a strategy is the most efficient way of focusing the language model’s attention.
We consider clustering algorithms for discovering small groups of related code; we call these tuples. This extends REGAL [3], which clusters programs solving similar problems by assuming each program is paired with a natural language description of the problem it solves, and clustering embeddings of those descriptions. But programs solving similar problems do not necessarily have similar structure. We therefore instead first summarize the code source itself by prompting a language model, then cluster based on the similarity of these new descriptions.
Once identified, each tuple is used in two stages. The first is to retrieve relevant already-abstracted functions from the Librarian library: for a given tuple of programs, relevant functions are retrieved by prompting a language model with the entire existing existing library and the original input programs, then instructing it to identify which functions would be useful. The retrieved library functions and the original programs in the tuple are then provided as critical context to the language model to propose a sample budget of $K$ candidate refactorings.
4.2 Rank with compression
Once sampled, all $K$ candidate refactorings are passed through a sort-and-filter evaluation harness to select which one scores the highest on refactor quality and that maintains (or improves) test accuracy compared to the original. If no such candidate exists, the original code is preserved, maintaining existing functionality.
New library functions in the selected refactor are saved into the Librarian library for potential use in downstream refactoring of other programs. We provide the full algorithm in Appendix A.
5 MiniCode
MiniCode evaluates a code agent’s capability to identify abstractions across implementations and design reusable libraries. In order to measure these capabilities, our benchmark presents agents with a collection of code sources, then asks agents to refactor the code sources into a unified library alongside refactorings of the original code sources. There are two key desiderata for collections of code sources: The collections must be compressible, in that there exists a latent shared library abstraction, and verifiable, so that we can measure how well refactored code sources preserve functional correctness. We source problems from two domains: Competition coding and synthesized repositories (Table 1).
Agents are expected to interact with MiniCode via the terminal. We structure the benchmark as refactoring a multi-package Python repository, where each code source in a collection is a Python package in a subdirectory. This requires knowledge of basic bash commands for exploring repositories, editing code, and running tests, as well as how to manage complex, multi-package Python libraries.
CodeContests
Competition problems are crafted with specific variations of algorithmic approaches in mind, resulting in both shared latent concepts and required test cases. As a result, competition coding is naturally both compressible and verifiable.
Each collection consists of multiple code sources, each containing a solution to a competition programming prompt, with associated tests for verification. We take solutions, prompts, and tests from CodeContests [19], a dataset consisting of competition coding problems. Each code source in the collection is structured as a subdirectory consisting of the task description in PROBLEM.md, the initial solution in main.py, and a script to run tests in run.sh. Agents are instructed to create a library.py file, which should is imported into each code source. Since CodeContests has no external dependencies on Python packages, this can be done without explicit structuring as a Python package.
Repositories
For the second domain of synthesized repositories, we propose a data-generating process that first produces project ideas, then generates variations of those project ideas tailored to specialized use-cases. This allows us to control the complexity and degree of overlap between each code source in a collection. Each code source in the repositories domain is comprised of a task description, source code, and test cases for functionality and correctness. MiniCode includes both small repositories (approximately 200 lines of source code each) and large repositories (approximately 6.5k lines of source code each), both of which represent realistic settings where different people with different needs use language models to help them write software for their particular use cases. The refactoring agent is tasked with extracting re-usable functions from across code repositories, and re-writing the original code source repositories to use them.
We approximate the true distribution of code repositories, $p(\rho)$, with a generative process using latent textual library and repository descriptions. The entire space of code repositories is massive. To collect and group code repositories that share meaningful structure that can be abstracted into a useful code library, we (1) sample a textual library description; (2) sample use cases; and last (3) their programmatic implementations. This generative process naturally produces repository collections primed for refactoring. We sample from language models for each step of this process. Prompts are shared in Appendix E.
Each collection consists of multiple synthetic repositories as code sources, where each code source is a subpackage. Collections are obtained by transforming the original repository code sources into a multi-package Python library. Each source’s source code is extracted into a subpackage directory, and its tests into a corresponding test subdirectory. Agents are instructed to write a shared library in a shared subpackage named common. This common shared library must be imported and used to refactor each of the original code sources.
Table 1: MiniCode Statisics
Domain Sources Collections Avg LoC Avg Tests Gen by
Code Contests 300 30 87 10 Humans
Small Repositories 262 22 209 12 o4-mini
Large Repositories 20 10 6,433 101 Claude-Sonnet 3.7
6 Experimental Setup
Grouping Programs into Collections
To facilitate parallel application of Librarian and manage the dataset scale, we assume that semantically distant code sources will have minimal overlap in their optimal library functions. Therefore, our overall approach partitions the dataset into disjoint collections through clustering.
For CodeContests, these collections are constructed from an initial corpus of $\sim$9k problems with Python solutions: We first filter these code sources, removing those whose selected canonical solution is under 10 lines (minimal refactoring potential). For the remaining 4596 solutions we use a language model to generate textual descriptions of canonical solutions—emphasizing reusable components—which are embedded using OpenAI’s text-embedding-ada-002.
Agglomerative Clustering [20] is subsequently applied to these embeddings to partition the code sources into a predefined number of initial clusters, in our case 120. To create uniformly sized experimental units, we subsample each such cluster to form collections of 30 code sources. This collection size was empirically chosen because it balanced between the runtime of Librarian without limiting compression. We select 10 collections that we then use to evaluate our methods.
Code repositories are generated as disjoint collections through the generative process, which first samples project ideas then variations. For small repositories, we up sample $10-15$ code sources then form smaller collections of size $3$-$5$ code sources that exhibit the most intra-cluster similarity according to clustering using the same embedding and clustering technique. For big repositories, we sample two code sources for each of the $10$ collections.
REGAL Baselines
To evaluate the ability of our libraries to support reuse on new problems, we turn to the program synthesis tasks used in REGAL, where learned libraries are added to help the program synthesizer. We evaluate on the two domains published by the authors, Logo and Date. Because our clustering is inspired by REGAL but adds additional complexity, for fair comparison, we keep their setup the same and only augment the training using sample + MDL rerank procedure described in Section 4.1.
Code Contests
To evaluate Librarian on refactoring Code Contests we select 6 collections of 30 code sources (problems). In each collection we group the problems into tuples of size $3$. We set the sample budget to be $K=8$, since our ablations show that with larger $K$ we discover better libraries 3. We use the MDL objective for rankings.
The model used for sampling is OpenAI’s o4-mini [21]. To obtain MDL scores we use Qwen 2.5 7B Instruct [22] as a balance between quality, speed, and cost.
Code Agents
To fairly evaluate performance on the task by state-of-the-art systems, we use coding agents that advertise long-context ability to reason about, write, and refactor code repositories. Specifically, we use Claude Code (Cl) [23] which uses the Claude 3.7 Sonnet model, and OpenAI’s Codex (Cx) [24] which uses o4-mini.
All steps of the initial code sources for small code repositories are done by OpenAI Codex. We empirically observed that Claude Code has an affinity for generating large code repositories– the large code repositories are all generated by Claude Code.
We test whether code agents can refactor collections of code sources autonomously, without human intervention. Refactoring repositories with code agents involves planning and iterative (re-)implementation and testing. Code agents are prompted to perform each of these steps, with feedback from the unit tests. Agents must run and repair unit tests autonomously. We run coding agents multiple times per collection, logging their progress in checklists stored in text files. Our naming convention for agents is “Planner-Executor”. For example, “Cl-Cx” uses the Claude Code agent for planning and OpenAI Codex to implement the plan.
As before we use Qwen 2.5 7B Instruct [22] to obtain MDL scores.
@@ 6 Experimental Setup
Percentage Single Use Functions 38.03% ±4.88
Figure 2: Refactoring results for Librarian (w/ $K=8$) averaged over 10 Code Contests collections.
7 Results
In this section, we present the compression and correctness results with Librarian and agent baselines on MiniCode.
MiniCode-CodeContests On CodeContests, Librarian achieves a high final pass rate of 90.67% and significantly improves correctness, with pass rates increasing by 6.33% compared to the original code sources (Table 2). The method yields substantial compression: the refactored code, including the new library, shows an MDL ratio of 0.53 (a 47% reduction in MDL relative to the original). On average, Librarian generates libraries containing approximately 11 functions. These functions demonstrate good reuse, being called by around 5.2 programs on average, although 38.03% of them are used only once within their specific collection context.
Code agents fail to achieve both high correctness and compression on MiniCode-CodeContests. Across collections, the Codex agent achieves an average MDL ratio of 0.83, but a pass rate of 74.16%, much lower than Librarian’s rate of 90.67%. Similarly, the Claude agent reaches a higher pass rate of 82.50%, which is still lower than Librarian, but an MDL ratio of 1.04 which is more complex than the original collection.
MiniCode-Repositories Coding agents results on both small and large repositories are given in Table 2. During experiments, we found that the Claude Code agent provided superior plans for refactoring, while both the Claude Code and Codex agents performed satisfactorily at implementing a plan.
Despite using code agents with a human in the loop, existing code agents fail to produce effective refactorings. Often, the resulting refactored codebase bloats in size and complexity, particularly when using Claude Code for both planning and implementation. Given the poor performance of code agents on these repository-scale refactorings, we do not run the full sample-and-rerank pipeline. Supporting qualitative analysis in Section 8.4 discusses common failure modes of code agents on this task in repository-size refactorings.
Collection Agent MDLPass%
datapipe orig 1.0 1.0
Cl-Cl 9.1 1.0
Cl-Cx 2.2 1.0
state_machine orig 1.0 1.0
Cl-Cl 6.3 0.9
Cl-Cx 1.9 1.0
config_schema orig 1.0 1.0
Cl-Cl 2.5 0
Cl-Cx 2.1 1.0
cli_tools orig 1.0 1.0
Cl-Cl 4.0 1.0
Cl-Cx 3.1 0
cli_form orig 1.0 1.0
Cl-Cl 3.8 1.0
Cl-Cx 2.9 1.0
Collection MDLPass%
filesys_analyzer orig 1.0 1.0
Cl-Cl 1.6 1.0
query_language orig 1.0 1.0
Cl-Cl 1.4 1.0
knowledge_store orig 1.0 1.0
Cl-Cl 1.4 1.0
vm_emulator orig 1.0 1.0
Cl-Cl 1.6 1.0
finance_tracker orig 1.0 1.0
Cl-Cl 1.8 0.6
in_memory_db orig 1.0 1.0
Cl-Cl 1.5 1.0
text_editor orig 1.0 1.0
Cl-Cl 1.2 1.0
Table 2: The correctness (Pass %) and compression (MDL ratio %) of the original and refactored code sources for the agent baselines on a subset of the small and large repository collections in MiniCode. Libraries with syntax or runtime errors receive a pass rate of 0.
8 Analysis
We further analyze parts of Librarian, consider alternative objective functions, analyze the use of clustering while constructing MiniCode-CodeContests, and common failure modes of code agents.
8.1 Can these libraries be reused to solve downstream programming problems?
To understand the practical value of our library extraction work for program synthesis, we augment the existing library learning algorithm, REGAL with sampling and MDL-reranking. In REGAL, the learned library is used to solve downstream holdout program synthesis problems, and a simpler clustering algorithm is used. For fair comparison with REGAL we hold the clustering algorithm constant (using the simpler REGAL approach), but sample $K=5$ refactorings/libraries and rerank by MDL. The resulting system outperforms REGAL on holdout programming problems (Table 3).
These problems however are highly simple, which is representative of classic library learning work: Either drawing simple geometric designs (dataset Logo) or manipulating textual representations of dates (dataset Date). This helped motivate us to consider refactoring whole repositories.
Table 3: Solving downstream program synthesis tasks using learned libraries
@@ 8.1 Can these libraries be reused to solve downstream programming problems?
8.2 What objective function should a library learner use?
We run our method on CodeContests using MDL, number of tokens, and cyclomatic complexity [25] [^2] Cyclomatic complexity (CC) is a classic metric from the software engineering community that analyses the number of paths through a program’s control flow, and unlike MDL is only loosely coupled with program size. as objective functions on 6 collections (Figure 3) While minimizing MDL also minimizes the other two objectives, the converse is not true. This suggests that MDL is a pareto-optimal loss among the three objectives in this experiment.
To confirm that the library does indeed expose shared abstractions, we calculate the average number of times that each library routine is used. Scaling the inference budget to $K=8$ discovers better libraries, reusing each library function on average about 5 times.
In Appendix A.1, we find that MDL is also the objective that best correlates with human preferences in a small-scale human study.
Figure 3: Comparing 3 different objective functions for refactoring (different columns) according to different downstream success metrics (different rows), as a function of refactoring budget (horizontal axes).The values are averaged over 6 collections of CodeContests problems. Row 1: Optimizing perplexity also incidentally optimizes cyclomatic complexity and token count, but that the converse is not true. Row 2: refactored programs pass more test cases, even more than the original code itself. Row 3: increasing the refactoring budget results in more reusable library subroutines (such subroutines are called more times on average). Filtered/Raw: Using/Not Using tests to filter samples.
8.3 Clustering Analysis: CodeContests
We analyze the coherence of the clusters underlying collections in MiniCode-CodeContests. In particular, we compare clustering based on o4-mini generated code source descriptions against task descriptions. Since task descriptions in competition coding problems are designed to hide the algorithmic approach needed to solve problem, we expect that clusters based on code source descriptions are more coherent.
We use two measures to evaluate the thematic coherence of collections: Good collections should group code sources with a (1) concentrated and (2) identifiable set of shared conceptual tags, which for CodeContests are provided as ground truth (trees, graphs, etc.).
Normalized Tag Instance Entropy
This measures the concentration of tag instances within a collection, and is given by the entropy of the tag distribution for a given collection, normalized by the number of distinct tags in that collection. A lower normalized tag instance entropy (closer to 0) indicates higher thematic purity, meaning a small number of tag types are most prevalent.
Herfindahl-Hirschman Index (HHI) for Problem Presence
This measures tag concentration across distinct code sources in a collection. A higher HHI signifies that the problems are collectively characterized by a smaller, more focused set of tags.
We provide the full formal definitions of both measures in Appendix B.
Figure 4 shows our clustering approach yields more thematically coherent clusters, evidenced by achieving lower entropy and higher HHI values across the entire tested range of $N$.
8.4 Common Failure Modes of Coding Agents
Failing to follow the plan.
Refactoring large repositories requires editing code over a long time horizon. We instruct agents to plan hierarchically by first recording their plan in natural language, then implementing that plan. However, agents often ignore the plan, particularly with Codex as the implementation agent, potentially refusing to edit the code at all. We hypothesize that this behaviour is caused by the refactoring task itself – the original code source passes all tests, and agents are trained to prioritize correctness over compression.
Failing to use the library.
When the implementation agent does follow the plan, it often fails to do so effectively. Claude Code, for example, will write a library, import it to the proper place, then rewrite the library rather than use it. We provide further examples in Appendix G.

### v3

@@ 5 Librarian: Refactoring Code to Create Libraries
$\mathcal{L}_{c},\left\{\rho^{\prime}_{c,i}\right\}_{i=1}^{|c|}$ $=\operatorname*{arg\,min}_{\mathcal{L},\left\{\rho^{\prime}_{c,i}\right\}_{i=1}^{|c|}\in\textsc{Sample}_{K}\left(c\right)}\ell\left(\mathcal{L},\left\{\rho^{\prime}_{c,i}\right\}_{i=1}^{|c|}\right)\text{, for each }c\in\textsc{Cluster}_{S}(\left\{\rho_{n}\right\})$ (5)
The approach above ignores the fact that library abstractions discovered in one cluster might be useful in another cluster. A more sophisticated approach accumulates a library across clusters, and when refactoring a cluster, adds the accumulated library to the prompt. This lets abstractions discovered earlier carry forward across the collection. Appendix˜A describes this extension.
6 What Makes a Good Refactoring?
We compare different metrics $M$ measuring the quality of a refactoring:
Tokens measures the total number of tokens in the refactored files and in the library. It minimizes program size, but not at the expense of creating a bloated library: Simply replacing every program with its own monolithic library function would not improve the tokens metric, because it measures library size as well. Concretely, $M_{\text{tokens}}(\mathcal{L},\{\rho^{\prime}_{n}\})=\textsc{Tokens}(\mathcal{L})+\sum_{n}\textsc{Tokens}(\rho^{\prime}_{n})$.
Minimum Description Length (MDL) evaluates the negative log probability under a reference distribution, taking into account both the library and refactored sources. Concretely, $M_{\text{MDL}}(\mathcal{L},\{\rho^{\prime}_{n}\})=-\log p_{\text{LM}}(\mathcal{L})+\sum_{n}-\log p_{\text{LM}}(\rho^{\prime}_{n}\mid\mathcal{L})$, where $p_{LM}(\rho_{n}^{\prime}|\mathcal{L})$ is concatenating the library and the program into one prompt, but only counting the perplexity of the later program tokens. This has a Bayesian justification: The MDL library is the maximum aposteriori estimate of $\mathcal{L}$ given conditionally-independent programs. We use Qwen-2.5-3B as our reference language model, as it is modern, open, and has publicly-available endpoints for querying logits, which is required for scoring refactorings. To confirm that our MDL optimization results are not model specific, we computed MDL values for 15 Code Contests clusters, each of 50 valid refactoring candidates, using Qwen and Llama-3.2-3B models. We found a 94% agreement in the minimum MDL candidates from both models.
Cyclomatic Complexity (CC) is a longstanding metric from the software engineering community which measures the number of linearly independent paths through a program’s control flow graph (McCabe, 1976). Smaller programs often have lower cyclomatic complexity. It is equivalent to defining $M_{\text{CC}}(\mathcal{L},\{\rho^{\prime}_{n}\})=\textsc{cc}(\mathcal{L})+\sum_{n}\textsc{cc}(\rho^{\prime}_{n})$ and $\textsc{cc}(\rho)=E(\rho)-N(\rho)+2P(\rho)$, where $E$, NN, and $P$ measure the number of control flow edges, nodes, and connected components, respectively.
Maintainability Index (MI) is a modern software engineering metric combining several other metrics, including lines-of-code, cyclomatic complexity, and Halstead volume into a single score. Higher MI values are intended to indicate easier-to-maintain code, so we define $M_{\text{MI}}(\mathcal{L},\{\rho^{\prime}_{n}\})=-\textsc{MI}(\mathcal{L})+\sum_{n}-\textsc{MI}(\rho^{\prime}_{n})$.
6.1 Asymptotic behavior of metrics in large-sample regime
Are these metrics equally effective at encouraging modular and reusable libraries? To answer this question, we run Librarian on 15 CodeContests (each of three files) using MDL, tokens, maintainability index, and cyclomatic complexity, while varying the inference-time sample budget $K$ (Figure˜2,7). We use Best@k estimator for the expected value of metrics for all $k\leq K$ (we describe the estimator and prove its correctness in Appendix F). Tokens and MDL separate cleanly from classic software engineering metrics: Optimizing tokens/MDL, both of which essentially compress the original programs, does not yield steady improvements in MI/CC, and vice-versa. To understand whether these libraries expose shared abstractions, we examine the average number of times that each library routine is used, and the average number of library invocations per library function. This teases apart tokens and MDL: Optimizing MDL yields more more reusable libraries (used about $8\times$ per task), with each function called more often (called about $2.2\times$ per function)—exceeding the other metrics we consider. See Appendix Figure˜7 for full results, including Cyclomatic Complexity.
Figure 2: (A) Asymptotic behavior of metrics for scoring libraries and refactorings (columns) varying refactoring budget (horizontal axes). (B) Comparing metrics via proxies of downstream library quality (total library usage and average calls per library function), for which MDL>>Tokens>>MI. All results are estimated using Best@k. See also Appendix Figure˜7.
Figure 3: Best@K MDL ratio. Increasing sample budget improves MDL on Transformers.
Studying these metrics at large $k$ allows understanding their inference-time scaling behavior. While the underlying metric itself improves (with diminishing returns), this was not to be taken for granted: The backend language model must produce sufficient diversity to steadily improve these metrics. Prior state-of-the-art, such as Stengel-Eskin et al. (2024), instead take a single sample, but the results here suggest benefit from further test-time search, and indeed, real-world repos benefit from steady improvement with increased samples (Figure 3). But our proxies of library utility plateau much earlier, around $k=20$ samples, suggesting large $k$ is unnecessary in practice: Effective library building benefits from test-time compute, but does not demand an exorbitant amount of it.
6.2 Human evaluation of refactoring metrics
Figure 4: Human evaluation of different refactoring objectives. Judges compare pairs of refactorings that both pass all test cases. MDL aligns best with human preferences.
We perform a human study to corroborate the findings of Section˜6.1 using the exact same CodeContests clusters (Appendix˜E). The human study compares tokens, MDL, and Maintainability Index by (1) refactoring clusters into libraries, (2) presenting human participants with the original sources and their refactorings under pairs of metrics, and (3) eliciting pairwise preferences from human participants.
Humans prefer MDL-minimizing libraries (Figure˜4), and although the preference is only statistically significant for MDL vs. MI, the data suggest a rank-order preference of MDL>Tokens>MI; Tokens vs. MI was slightly below the threshold of significance. We ran $N=14$ participants (eliciting 129 judgments), and although more participants would make more of these comparisons statistically significant, already we see a general preference for compression-based metrics (MDL and Tokens) with only MDL crossing the threshold of statistical significance. Indeed, minimizing MDL is not quite the same as minimizing tokens, as Figure˜5 illustrates: We believe basically every human coder would prefer the MDL-minimizing program in this example.
We therefore adopt $M_{MDL}$ as the primary objective in the remainder of this paper: In addition to support from this human study, (1) Bayesian arguments support MDL; (2) corner cases in the style of ‘Perl golf’ provide existence proofs of the liability of merely minimizing tokens; and (3) reasonable proxies for library reuse favor MDL (Section˜6.1).
Figure 5: Example where tokens and MDL diverge: Obfuscating the original library definitions (left) by shortening variable names (right) reduces tokens but increases MDL.
7 What we learn from running Librarian on MiniCode
We empirically study Librarian on MiniCode with the goal of understanding (1) the degree to which library abstractions are reused across programs, (2) how our method compares to state-of-the-art library learning on existing datasets, and (3) whether Librarian holds value for real-world repos.
Librarian discovers reusable functions for competition programming–but some functions are only called once.
We test on CodeContests with a cluster size of $S=3$ and a sample budget of $K=8$ draws from o4-mini, as reasoning models perform well on competition programming. [^2] Code agents such as Codex, Claude Code, and others underperformed o4-mini (Appendix I) Table˜2 shows that the resulting refactors and libraries approximately halve the MDL, which incidentally reduces program size as well (44% relative reduction in token count). Pass rate modestly improves as an incidental consequence of sampling and filtering with test cases. Libraries average 10 functions, each heavily reused: Averaging 5 uses per function within tasks comprising only 10 programs. But almost 40% of library functions are only used once. Why is that?
A signature of the MDL objective is a preference for whatever a language model assigns high apriori probability to. Although a single-use function does not reduce line count or tokens —the function could simply be inlined—it improves MDL if it yields a more natural decomposition of the target programs. Indeed, human-written libraries sometimes include functions that are seldom used, provided they serve as a conceptually modular abstraction. We therefore see single-use functions as a feature, not a bug. See Appendix J for an example refactoring candidate on CodeContests.
Are these libraries useful for solving new, unseen programming problems?
Library learning has long sought to learn libraries from training programs which then help solve new unseen program synthesis tasks. The Logo and Date datasets fit within this paradigm. Recently REGAL improved the state-of-the-art on these library learning datasets. Because our clustering is heavily inspired by REGAL, for fair comparison, we keep exactly their clustering setup but add MDL-based reranking using $K=5$ samples. Despite the simplicity of these datasets, we find value in our more complicated method. Table˜3 shows that sampling and reranking by MDL yields up to a 41.8% relative improvement in solve rate on unseen programming problems, and that even when the gains are more modest, we still improve upon the state-of-the-art. But these are relatively simple problems solvable with about ten lines of code—does this work in the real world?
Table 2: Results for Librarian on 10 Code Contests tasks ($K=8,S=3$)
@@ 7 What we learn from running Librarian on MiniCode
% Single Use Functions 38.03% ±4.88
Table 3: Solving holdout test program synthesis tasks using learned libraries.
@@
Real-World Refactoring. The HuggingFace Transformers library is used by nearly 400k GitHub projects. We deploy Librarian to 10 source files, using Claude Code to sample $K=15$ refactorings per cluster of size $S=5$, believing an agent such as Claude Code would excel at repo-level edits. Librarian distilled repeated abstractions such as MLPs, Attention, Decoder classes, RoPE helper functions, etc., lowering MDL to 67.2% of its original value while still passing all integration tests. The top-3 refactorings based on MDL have an average of $18\pm 4.4$ abstractions (functions, classes) in the library, each of which is called on average $4.59\pm 0.39$ times in the refactored models. For Diffusers, scheduler clusters yielded top-3 MDL refactorings with an average of $12.3\pm 1.6$ functions and $3.0\pm 0.4$ calls per function, while UNet refactorings produced richer abstractions with an average of $17.0\pm 5.6$ functions/classes and $3.43\pm 0.67$ calls each.
Refactoring at scale proved expensive: Each refactoring took approximately 30 minutes to generate and test. But this is a one-off cost, and in our view, the refactored Transformers and Diffusers sources are much cleaner, and the new library is transparently reusable (Figure˜6). To the best of our knowledge, this is the first time any library learning algorithm has been successfully applied to real-world software projects.
Figure 6: Representative result for refactoring HuggingFace Transformers using Librarian
Learned libraries from these real-world codebases are useful for unseen downstream refactoring tasks. When a library learned on one cluster of Transformer files (5 models) is applied to refactor a second cluster, Librarian reduces the unseen cluster’s MDL to 73% of the its original value, with an average of 3.0 calls per library function. This demonstrates that Librarian learned libraries that can be repurposed to more compactly rewrite unseen real-world code sources.
Human performance. HuggingFace is concurrently refactoring the Transformers library. On the same 10 source files used above, their refactoring achieves an MDL compression ratio of 66.5%. To estimate an independent ceiling, we the authors refactor the task ourselves and obtain 62%. This suggests room for a little improvement beyond Librarian’s 67.2%, but not by much (from inspection, specifically with better implementations of mixture-of-experts and vision models).